# Emergence of a second law of thermodynamics in isolated quantum systems

### v2

@@ I Introduction
Irreversible evolution observed in nature is characterized by the second law of thermodynamics [1, 2, 3, 4], and is thus quantitatively captured by the increase in entropy over time of an isolated system. From the perspective of microscopic physics, however, all laws are time-reversal invariant and, thus, if a process is allowed by these laws, its reverse process must be, too. This means that entropy could, in principle, also decrease. This apparent conflict between the reversibility of physics on the microscopic level and the irreversibility of macroscopic physics, sometimes called Loschmidt’s paradox, has been debated since the earliest days of classical statistical mechanics [5, 6, 7, 8]. As for the entropy, for classical systems, Liouville’s theorem implies that the Gibbs entropy formula is constant under Hamiltonian evolution [9], and also for quantum systems, the unitary evolution generated by the Schrödinger equation preserves the von Neumann entropy. In this Letter, we address this issue from the perspective of quantum systems, determining the sense in which entropy can increase over time in isolated quantum systems.
@@
With regards to quantum mechanics, despite substantial theoretical arguments [21, 22, 23, 24, 25, 26, 27, 28] and experimental evidence [29, 30, 31, 29] for the equilibration of isolated quantum systems, no consensus exists about how isolated quantum systems equilibrate in the sense of something akin to the second law, as unitary evolution preserves the von Neumann entropy. By using the modern theory of equilibration of isolated quantum systems [22, 23] we show how the second law of thermodynamics is satisfied on average. In this framework, the system is examined relative to an arbitrary observable represented by a self-adjoint operator $O=\sum_{i=1}^{r}O_{i}\Pi_{i}^{O}$, where $r$ is the number of possible measurement outcomes. Assuming finite Hilbert space dimension $d$ (with $r<d$) and identifying the measurement outcomes with macrostates introduces, in a sense which we will quantify, the necessary lack of information, and ensuing emergent irreversibility.
@@ IV Discussion
While this work focused on the Shannon observable entropy, we have proved analogous statements for the observational entropy in Appendix II.2, showing that the condition $r/d_{\mathrm{eff}}\ll 1$ is not only sufficient for equilibration of the Shannon observable entropy but also for the observational entropy. Still, one can construct examples where the Shannon observable entropy fluctuates on the order of $\log r$ and thus does not equilibrate, whereas the observational entropy’s fluctuations are relatively small because of the additional Boltzmann term which is of order $\log d$. This could lead one to conclude that the observable equilibrates while in fact it does not, as the Shannon entropy shows. Regardless of its role in understanding equilibration, the observational entropy does quantify the ignorance resulting from multiple states being compatible with a given measurement outcome [52], which the Shannon observable entropy alone does not capture. As we have shown, this many-to-one relationship between states and outcomes instead dictates the dynamics of the Shannon entropy, depending on the dynamical participation of those states as encoded in $d_{\mathrm{eff}}$.
@@
Given that the ratio of observable outcomes to effective dimension in the case of a small subsystem of a much larger total system is generally favorable to equilibration (the former being bounded by the dimensionality of the subsystem), our results can likely be used to constrain the behavior of the von Neumann entropy of a subsystem. Since our focus was on the overall behavior of isolated systems, we have not addressed this here. Additionally, one may ask how our analysis generalizes to the context of other notions of entropy such as the family of Tsallis or Rényi entropies [53, 54, 55], by defining them relative to an observable in analogy to how we define the Shannon observable entropy. By combining entropy continuity bounds such as those of e.g. [54, 56] together with Lemma I.1 for PVMs or Lemma C for POVMs (see Appendix), respective formulation of the second law might be obtained for such entropies.
A natural question following from the results presented here is that of how our statement of the second law compares to the historical formulations of Clausius [57] (“No process is possible, the sole result of which is the transfer of heat from a cold body to a hot body”) or Kelvin-Planck [58, 2] (“No process is possible, the sole result of which is that a body is cooled and work is done”). A rigorous analysis of these statements would require a definition of work, heat and temperature for subsystems of isolated quantum systems, which would be relative to the observable the agent can use to examine and manipulate the system. For example, for the spin-system in our numerical analysis in Fig. 1, a suitable observable is the (intrinsic) total magnetization in a given direction of space, which might be seen as a thermodynamic state variable. Note that the formulations of both Clausius and Kelvin-Planck can be turned into statements about a decrease in thermodynamic entropy [2] of the systems, which, given compatibility of the thermodynamic entropy with the Shannon observable entropy, is forbidden by Theorem 2 up to stochastic fluctuations.
@@
Finally, our work corroborates the findings of previous studies noting that non-integrability does not appear to be a necessary requirement for equilibration relative to observables [35, 38, 37, 36]. As we showed in (10), a large $d_{\mathrm{eff}}/r$ suffices for equilibration of the observable probability to the equilibrium distribution, and thus for our formulation of the second law to hold, and this does not a priori require non-integrability. Whether the system is integrable or not may still affect some details like the speed of equilibration or tightness of the bounds as noted in [35, 24]. In other contexts as well, a connection between integrability and effective dimension has been hinted at [59, 60], which leads us to pose the question of whether $d_{\mathrm{eff}}$ could allow for a quantitative definition of integrability in relation to the equilibration of quantum systems.
@@ I.1 Proofs for averaged entropy differences
Theorem 2 (Second law on average).
$\left\langle\left|S_{\mathrm{Sh}}^{O}[\rho(t)]-S_{\mathrm{Sh}}^{O}[\omega]\right|\right\rangle_{T}\leq\log(1+r)\eta_{\varepsilon,T}+H_{2}[\eta_{\varepsilon,T}],$ (26)
@@
We subdivide the proof of this statement into two steps: (1) We express the left-hand side of Eq. (26) in terms of population differences, giving us a bound on the entropy difference on average expressed as a function of the average difference in populations between $\rho(t)$ and the infinite-time averaged state $\omega$. (2) Then, we use the result from Lemma I.1, which we can then insert into the bound found in the first step, finalizing the proof of this Proposition.
@@ I.2 Proofs for fluctuation theorems
$E[X]=\frac{\log(1+r)}{2}\sqrt{\frac{r}{d_{\mathrm{eff}}}}+H_{2}\left[\frac{1}{2}\sqrt{\frac{r}{d_{\mathrm{eff}}}}\right]=:\delta,$ (38)
as a direct consequence of Proposition 2. Setting $\kappa:=\sqrt{\delta}$ yields
@@ II.1 A second law for the von Neumann entropy?
II.2 Generalization for the observational entropy

### v3

@@ I Introduction
Irreversible evolution observed in nature is characterized by the second law of thermodynamics [1, 2, 3, 4], and is thus quantitatively captured by the increase in entropy over time of an isolated system. From the perspective of microscopic physics, however, all laws are time-reversal invariant and, thus, if a process is allowed by these laws, its reverse process must be, too. This means that entropy could, in principle, also decrease. This apparent conflict between the reversibility of physics on the microscopic level and the irreversibility of macroscopic physics, sometimes called Loschmidt’s paradox, has been debated since the earliest days of classical statistical mechanics [5, 6, 7, 8]. As for the entropy, for classical systems, Liouville’s theorem implies that the Gibbs entropy formula is constant under Hamiltonian evolution [9], and also for quantum systems, the unitary evolution generated by the Schrödinger equation preserves the von Neumann entropy. In this article, we address this issue from the perspective of quantum systems, determining the sense in which entropy can increase over time in isolated quantum systems.
@@
With regards to quantum mechanics, despite substantial theoretical arguments [21, 22, 23, 24, 25, 26, 27, 28, 29, 30] and experimental evidence [31, 32, 33, 31] for the equilibration of isolated quantum systems, no consensus exists about how isolated quantum systems equilibrate in the sense of something akin to the second law, as unitary evolution preserves the von Neumann entropy. Progress has been made in showing that, for example, the diagonal entropy of closed, driven (i.e. energy non-conserving) quantum systems at some time $t>0$ is non-decreasing from its value at $t=0$ [22, 23], or that the observational entropy of most states in the Hilbert space is close to the maximum [30]. However, a proof that the entropy of an isolated system tends to an equilibrium value over time, and a general exposition of the conditions under which this occurs, has been missing. We fill this gap using the modern theory of equilibration of isolated quantum systems [24, 25] showing how the second law of thermodynamics is satisfied on average.
In this framework, the system is examined relative to an arbitrary observable represented by a self-adjoint operator $O=\sum_{i=1}^{r}O_{i}\Pi_{i}^{O}$, where $r$ is the number of possible measurement outcomes. Assuming finite Hilbert space dimension $d$ (with $r<d$) and identifying the measurement outcomes with macrostates introduces, in a sense which we will quantify, the necessary lack of information, and ensuing emergent irreversibility.
@@
Unlike the von Neumann entropy, the Shannon observable entropy can change in time for isolated systems. Indeed, we find that over the course of the evolution of an isolated system, the Shannon observable entropy is close to that of the infinite-time-averaged state
@@
also known as the equilibrium state in the context of equilibration on average:
@@ IV Discussion
While this work focused on the Shannon observable entropy, we have proved analogous statements for the observational entropy in Appendix II.2, showing that the condition $r/d_{\mathrm{eff}}\ll 1$ is not only sufficient for equilibration of the Shannon observable entropy but also for the observational entropy. Still, one can construct examples where the Shannon observable entropy fluctuates on the order of $\log r$ and thus does not equilibrate, whereas the observational entropy’s fluctuations are relatively small because of the additional Boltzmann term which is of order $\log d$. This could lead one to conclude that the observable equilibrates while in fact it does not, as the Shannon observable entropy shows. Regardless of its role in understanding equilibration, the observational entropy does quantify the ignorance resulting from multiple states being compatible with a given measurement outcome [54], which the Shannon observable entropy alone does not capture. As we have shown, this many-to-one relationship between states and outcomes instead dictates the dynamics of the Shannon entropy, depending on the dynamical participation of those states as encoded in $d_{\mathrm{eff}}$.
@@
Given that the ratio of observable outcomes to effective dimension in the case of a small subsystem of a much larger total system is generally favorable to equilibration (the former being bounded by the dimensionality of the subsystem), our results can likely be used to constrain the behavior of the von Neumann entropy of a subsystem. Since our focus was on the overall behavior of isolated systems, we have not addressed this here. Additionally, one may ask how our analysis extends to other notions of entropy. One such example is the diagonal entropy defined relative to the energy eigenbasis as in [22, 23], which is trivially constant for an isolated quantum system. When defining a diagonal entropy in the context of observables akin to [55], however, a non-trivial behavior is recovered (details in Appendix II.3). Other examples include the family of Tsallis or Rényi entropies [56, 57, 58] defined relative to an observable in analogy to how we define the Shannon observable entropy. By combining entropy continuity bounds such as those of e.g. [57, 59] together with Lemma I.1 for PVMs or Lemma C for POVMs (see Appendix), respective formulation of the second law might be obtained for such entropies.
On the classical second law formulations.
A natural question following from the results presented here is that of how our statement of the second law compares to the historical formulations of Clausius [60] (“No process is possible, the sole result of which is the transfer of heat from a cold body to a hot body”) or Kelvin-Planck [61, 2] (“No process is possible, the sole result of which is that a body is cooled and work is done”). A rigorous analysis of these statements would require a definition of work, heat and temperature for subsystems of isolated quantum systems, which would be relative to the observable the agent can use to examine and manipulate the system. In certain special cases of equilibrating systems, the classical thermodynamic notions of entropy can be recovered – in which case we have thermalization. One paradigmatic setting for thermalization comprises a subsystem $S$ weakly coupled to a larger environment $E$. Under suitable assumptions as exemplarily outlined in the seminal works [62, 50, 35], the equilibrium state $\omega$ on the global system is typically well approximated by the thermal state $e^{-\beta H_{S}}/Z_{S}$ of $S$. Here, $\beta$ is some temperature parameter associated with the energy and $Z_{S}$ the canonical partition function ensuring normalization of the thermal state. When the observable under consideration is the Hamiltonian $O=H_{S}\otimes\mathds{1}_{E}$ of the subsystem, the equilibrium Shannon observable entropy is given by the Gibbs thermal entropy $S_{\rm Obs}^{H_{S}}[\omega]\approx\log Z_{S}+\beta\langle H_{S}\rangle$ with Boltzmann constant $k_{B}=1$, up to deviations coming from the difference between the actual equilibrium state and the Gibbs state, as we outline in more detail in Appendix II.4. Theorem 1 then states that as the subsystem’s state evolves, its Shannon observable entropy (relative to the subsystem Hamiltonian) also converges to the thermal equilibrium value, in agreement with classical formulations of the second law. How a Gibbs-like form of the entropy relative to an observable can also be recovered in a more general context beyond the system-environment paradigm was investigated in [55], finding that such a form can encode equilibrium behaviour in various examples.
@@
Finally, our work corroborates the findings of previous studies noting that non-integrability is not a necessary requirement for equilibration relative to observables [37, 40, 39, 38]. As we showed in (11), a large $d_{\mathrm{eff}}/r$ suffices for equilibration of the observable probability to the equilibrium distribution, and thus for our formulation of the second law to hold, and this does not a priori require non-integrability. Whether the system is integrable or not may still affect some details like the speed of equilibration or tightness of the bounds as noted in [37, 26]. In other contexts as well, a connection between integrability and effective dimension has been hinted at [63, 64], which leads us to pose the question of whether $d_{\mathrm{eff}}$ could allow for a quantitative definition of integrability in relation to the equilibration of quantum systems.
@@ I.1 Proofs for averaged entropy differences
See 1
@@
We subdivide the proof of this statement into two steps: (1) We express the left-hand side of Eq. (12) in terms of population differences, giving us a bound on the entropy difference on average expressed as a function of the average difference in populations between $\rho(t)$ and the infinite-time averaged state $\omega$. (2) Then, we use the result from Lemma I.1, which we can then insert into the bound found in the first step, finalizing the proof of this Theorem.
@@ I.2 Proofs for fluctuation theorems
$E[X]=\frac{{\log(r-1)}}{2}\sqrt{\frac{r}{d_{\mathrm{eff}}}}+H_{2}\left[\frac{1}{2}\sqrt{\frac{r}{d_{\mathrm{eff}}}}\right]=:\delta,$ (38)
as a direct consequence of Theorem 1. Setting $\kappa:=\sqrt{\delta}$ yields
@@ II.1 A second law for the von Neumann entropy?
II.2 Extension for the observational entropy
@@
II.3 On the diagonal entropy
In the context of the diagonal entropy as for example used in the Refs. [22, 23], a partial extension of our results is also possible. The mentioned references define the diagonal entropy relative to the orthonormal energy eigenbasis $\{\left|\right.\!{\varepsilon_{n}}\!\left.\right\rangle\}_{1\leq n\leq d}$ of the system’s Hamiltonian $H$ assumed to be non-degenerate. For a quantum state $\rho$ the diagonal entropy is then given by
$S_{\rm d}[\rho]=-\sum_{n=1}^{d}\rho_{nn}\log\rho_{nn},$ (72)
where $\rho_{nn}=\langle{\varepsilon_{n}}|{\rho|\varepsilon_{n}}\rangle$ are the diagonal entries of $\rho$ relative to the energy eigenbasis. In case of an isolated quantum system, however, the Hamiltonian is time-independent, and consequently also the diagonals $\rho_{nn}(t)$ of the state $\rho(t)=e^{-iHt}\rho(0)e^{iHt}$. The diagonal entropy relative to the Hamiltonian $S_{\rm d}[\rho(t)]$ thus trivializes to a constant and does not quantify whether and how an isolated quantum system equilibrates.
When defining a diagonal entropy more generally relative to some orthonormal basis =$\mathcal{B}=\{\left|\right.\!{n}\!\left.\right\rangle\}_{1\leq n\leq d}$ of the Hilbert space of interest allows for more informative conclusions. The diagonal elements $\rho_{nn}=\langle{n}|{\rho|n}\rangle$ of $\rho$ can be taken in the basis representation of the states $\left|\right.\!{n}\!\left.\right\rangle$ instead of $\left|\right.\!{\varepsilon_{n}}\!\left.\right\rangle$. A diagonal entropy $S_{\rm d}^{\mathcal{B}}$ relative to $\mathcal{B}$ can then be defined akin to the diagonal entropy in Eq. (72). Since the diagonal vector of $\rho$ is in general $d$-dimensional, Lemma I.1 does not directly constrain the diagonals because in this case $r=d$ does not satisfy $r\ll d_{\rm eff}$. However, when considering an observable $O$ with $r$ outcomes satisfying $r\ll d_{\rm eff}$ as in the main text, a preferred basis can be defined as in [55]. Relative to the observable basis, the diagonal entropy can satisfy a second law in a sense similar to Theorem 1. Decomposing the eigenprojectors of the observable,
$\Pi_{i}^{O}=\sum_{k=1}^{V_{i}}|{i,k}\rangle\!\langle{i,k}|,$ (73)
into a sum of $V_{i}=\textnormal{Tr}[\Pi_{i}^{O}]$ 1-dimensional orthogonal projectors $|{i,k}\rangle\!\langle{i,k}|$ defines a basis $\mathcal{B}_{O}=\{\left|\right.\!{i,k}\!\left.\right\rangle\}_{i,k}$. Within each subspace $1\leq i\leq r$ the choice of basis is not uniquely defined by $O$, and arbitrary unitary transformations within the subspace $\{\left|\right.\!{i,k}\!\left.\right\rangle\}_{k}$ for fixed $i$ map the basis to another valid orthonormal one. Chosing the basis such that the diagonal entropy $S_{\rm d}[\rho]$ is maximized yields matrix elements,
$\langle{i,k}|{\rho|i,k}\rangle=\langle{i,k^{\prime}}|{\rho|i,k^{\prime}}\rangle=\frac{p_{i}}{V_{i}},$ (74)
which are equal within each subspace defined by the observable’s eigenspace projectors $\Pi_{i}^{O}$. In this case, the diagonal entropy equals the observational entropy defined by Eq. (9), as shown in [55],
$\max_{\mathcal{B}_{O}}S_{\rm d}[\rho]$ $=-\min_{\mathcal{B}_{O}}\sum_{i=1}^{r}\sum_{k=1}^{V_{i}}\frac{p_{i}}{V_{i}}\log\frac{p_{i}}{V_{i}}$ (75)
$=S_{\rm Obs}^{O}[\rho].$ (76)
As we have shown in Appendix II.2, this entropy also satisfies a second law bound like the Shannon observable entropy and thus also the diagonal entropy defined in the context of sufficiently coarse-graining observable.
II.4 Special case of thermodynamic entropy
When comparing our formulation of the second law based on Theorem 1 with the historical ones by Clausius [60] or Kelvin-Planck [61] certain additional assumptions have to be imposed on the setting under consideration. These classical versions of the second law are framed through the concept of thermalization. By contrast, the present work adopts the broader notion of equilibration where the focus is on how observables evolve toward stable equilibrium values. To recover the standard thermodynamic notions of entropy from our results, two separate concepts have to be introduced: (a) ‘General canonical typicality’ [62, 35] stating that most states of interest are close to a specific canonical state $\Omega_{S}$ in a sense which we specify in the following, and (b) the ‘thermal canonical principle’ stating that the canonical state $\Omega_{S}$ is close to the thermal state under appropriate assumptions [71, 50].
In this appendix, we argue how results on thermalization from previous works like Refs. [50, 62, 35] and standard techniques from statistical mechanics can be combined with our main Theorem 1 to recover the thermodynamic entropy as a special case of the Shannon observable entropy. To comprehensively arrive at this goal, we use the canonical ensemble where the isolated system is made up of two subsystems, one of which we call the system $S$ and the other the environment $E$. In the non-trivial case where $S$ and $E$ interact, $S$ is not isolated anymore, only the joint global system is.
Ultimately, we would wish to find that under reasonable assumptions which are to be specified, an initial state $\left|\right.\!{\psi}\!\left.\right\rangle$ on the global system yields a thermal equilibrium state on the system $S$. Being more precise, we wish to find that the equilibrium state $\omega_{\psi}=\langle|{\psi}\rangle\!\langle{\psi}|\rangle_{\infty}$ coming from $\left|\right.\!{\psi}\!\left.\right\rangle$, reduced on the system $S$ is close to the thermal Gibbs state at some inverse temperature $\beta$, meaning
$\textnormal{Tr}_{E}[\omega_{\psi}]\approx\tau_{S}=\frac{e^{-\beta H_{S}}}{Z_{S}},$ (77)
where $H_{S}$ is the system Hamiltonian and $Z_{S}=\textnormal{Tr}[e^{-\beta H_{S}}]$ is the partition function. Let us suppose for now that the relation (77) is satisfied; we return to this assumption later. Furthermore, we consider the system’s Hamiltonian $H_{S}$ as the observable of interest, $O=H_{S}\otimes\mathds{1}_{E}$. The Shannon entropy relative to $O$ evaluated at the equilibrium state $\omega_{\psi}$ then yields the usual thermodynamic entropy. By explicitly calculating
$p_{i}=\textnormal{Tr}\left[\left(\Pi_{i}^{H_{S}}\otimes\mathds{1}_{E}\right)\omega_{\psi}\right]=\frac{e^{-\beta\varepsilon_{i}}}{Z_{S}},$ (78)
where $\Pi_{i}^{H_{S}}$ is the projector on the $i$th energy eigenstate of $H_{S}$, the Shannon observable entropy can be determined to be
$S_{\rm Sh}^{H_{S}}[\omega_{\psi}]$ $=-\sum_{i=1}^{d_{S}}p_{i}\log p_{i}$ (79)
$=\log Z_{S}+\beta\langle H_{S}\rangle.$ (80)
Thus, assuming equality for (77), the Shannon observable entropy of the equilibrium state relative to the system’s Hamiltonian as the observable is in agreement with the thermodynamic entropy.
In practice, however, the relation (77) is only true approximately. Moreover, even though the deviations $\kappa:=\|\textnormal{Tr}_{E}[\omega_{\psi}]-\tau_{S}\|_{1}$ are small for most states, there are some where $\kappa$ significantly deviates from zero. It is also known that not all systems thermalize and thus quantifying these deviations has been the subject of extensive studies. In the remainder of this appendix, we provide pertinent references for how and under which assumptions the Gibbs state can be recovered. In our setting of interest, the total Hamiltonian is then given by
$H=H_{S}\otimes\mathds{1}_{E}+\mathds{1}_{S}\otimes H_{E}+H_{SE},$ (81)
where $H_{S}$ is the Hamiltonian of $S$ as before, and $H_{E}$ is the environment Hamiltonian $H_{E}$ and $H_{SE}$ are the interactions, assumed to be negligible compared to the other two terms. Conventionally, the initial state $\left|\right.\!{\psi}\!\left.\right\rangle$ is assumed to be a uniformly randomly picked state from a macroscopic slice $\mathcal{H}_{R}\subseteq\mathcal{H}_{S}\otimes\mathcal{H}_{E}$ from the full Hilbert space [62]. To recover the Gibbs state as in e.g. [50], the microcanonical slice is considered, which is spanned by all energy eigenstates of $H$ with energy eigenvalue within the window $[E,E+\Delta)$. Conventionally, it is assumed that $\Delta$ is small (on the scale of $E$ minus the ground state energy) but large enough that there is a macroscopic number of energy eigenstates within the window $E\pm\Delta$.
Starting with the first point (a), the canonical state $\Omega_{S}$ of the system $S$ is the one obtained by uniformly randomly picking a state from the macroscopic slice $\mathcal{H}_{R}$ and then reducing to $S$,
$\Omega_{S}=\textnormal{Tr}_{E}\left[\frac{\mathds{1}_{R}}{d_{R}}\right].$ (82)
The ‘general canonical principle’ as in Popescu et al. [62] then states that most states $\left|\right.\!{\psi}\!\left.\right\rangle\in\mathcal{H}_{R}$ will yield a reduced state on the system very close to the canonical state if the environment $E$ is sufficiently larger than the system $S$. Formally, the general canonical principle quantifies the average distance $\langle\|\textnormal{Tr}_{E}[|{\psi}\rangle\!\langle{\psi}|_{E}]-\Omega_{S}\|_{1}\rangle_{\psi}$ where the average $\langle\cdot\rangle_{\psi}$ is the Haar average over the subspace $\mathcal{H}_{R}$. This is however, not yet enough, as we wish to have that also the equilibrium state $\textnormal{Tr}_{E}[\omega_{\psi}]$ reduced on the subsystem is close to the canonical state. In the reference by Linden et al. [35], such a statement is shown, providing a bound for the average distance $\langle\|\textnormal{Tr}_{E}[\omega_{\psi}]-\Omega_{S}\|_{1}\rangle_{\psi}$. Note that there is a technical difference in how Popescu et al. and Linden et al. define the canonical state; while Popescu et al. define $\Omega_{S}$ according to eq. (82), Linden et al. define it as $\tilde{\Omega}_{S}=\langle\textnormal{Tr}_{E}[\omega_{\psi}]\rangle_{\psi}$. Due to linearity of the trace and the map $|{\psi}\rangle\!\langle{\psi}|\mapsto\omega_{\psi}$, both definitions agree.
The second point (b) is about the ‘thermal canonical principle’ and a statement about when canonical state $\Omega_{S}$ equals the Gibbs state $e^{-\beta H_{S}}/Z_{S}$ when we consider the microcanonical slice $\mathcal{H}_{R}$. This problem can be addressed using textbook methods from e.g. Ref. [71], where a standard assumption is that the interactions $H_{SE}$ are weak compared to the rest of the Hamiltonian. Such an assumption not needed for the more general canonical principle [62, 35]. In the reference [50], this special case is considered more rigorously by providing an exemplary derivation for how the microcanonical slice $\mathcal{H}_{R}$ yields the Gibbs state, given that the Hamiltonian spectrum is sufficiently dense (yielding a macroscopic slice $\mathcal{H}_{R}$).
In summary, we find that by combining results from the literature on typicality [62, 35] with thermal canonical typicality [71, 50] from statistical mechanics, the Shannon observable entropy of a suitably chosen (sub)system coincides with the standard thermodynamic notion of entropy from the canonical ensemble.